Suppose you decide to deposit $20,000 in a savings account that pays a nominal rate of 6%, but interest is compounded daily. Based on a 365-day year, how much would you have in the account after 6 months? (Hint: To calculate the number of days, divide the number of months by 12 and multiply by 365.)

A) $20,196.86

B) $21,639.49

C) $20,609.04

D) $19,990.77
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=20,000*(1+0.06/365)^(365*6/12)

=20,000*1.03045199

=$20609.04(Approx).